OpenStudy (king):

Help.......... L,M and N are midpoints on sides AB,BC,CA respectively.Area of triangle ABC is 48 units,then find area of triangle LMN......

OpenStudy (anonymous):

area of triangle LMN = \[1/2*(64\sqrt{3}/4)*\sqrt{3}/2 = 12\]

OpenStudy (king):

please explain fully i do not require only answer ........

OpenStudy (anonymous):

thats the right answer ...

OpenStudy (king):

yeah but i want the explanation as i already know the answer......

OpenStudy (anonymous):

The segment joining the midpoints of two sides of a triangle is parallel to the third side and half as long as the third side.

OpenStudy (king):

yeah i know that but wat do i do ahead?..

OpenStudy (anonymous):



OpenStudy (anonymous):

Since they're all midpointsegments, they're all half the length of the opposing side. What would that do to your area?

OpenStudy (anonymous):

Original area is 1/2*A*B Area of small triangle is 1/2* 1/2A * 1/2B and that is 1/4 of the original area, get it?

OpenStudy (king):

oh ohk thanx.......

OpenStudy (anonymous):

this triangle is divided into 4 equal parts so area of original triangle divided by 4